OpenStudy (anonymous):

Choose the correct simplification of the expression d3 • d5. d d15 d2 d8

OpenStudy (callisto):

Is the question \(d^3 \cdot d^5\) ?

OpenStudy (callisto):

Okay. \[x^a \cdot x^b = x^{a+b}\]Can you use this property to do your question now?

OpenStudy (anonymous):

is it d8

OpenStudy (callisto):

Yup :) Please type d^8 next time. This would be clearer :)

OpenStudy (anonymous):

k sorry

OpenStudy (callisto):

It's okay!! Don't worry!